OpenStudy (anonymous):

Help me show my work please? 3. Which equation represents a line with a slope -1/2 that passes through (-3, 4) Answer: 2y + x = 5

OpenStudy (vishweshshrimali5):

How did you approach the problem ?

OpenStudy (anonymous):

I tried to do it, and then asked my sister to help but, i'm confused.

OpenStudy (vishweshshrimali5):

Okay no problem

OpenStudy (vishweshshrimali5):

See when I am given the slope of a line and a point passing through it, I can use this form of the line: \(\large{(y-y_1) = m*(x-x_1)}\) Here, \((x_1,y_1)\) is the point lying on the line ; m is the slope of the line.

OpenStudy (vishweshshrimali5):

Now can you solve it ?

OpenStudy (anonymous):

hold on , let me try that

OpenStudy (vishweshshrimali5):

Go on.

OpenStudy (anonymous):

could you use my equation but, with different numbers with the form you gave me as an example, cause i still didn't get it right

OpenStudy (vishweshshrimali5):

No problem

OpenStudy (vishweshshrimali5):

I am going to change the question to this :\(\color{blue}{\text{Originally Posted by}}\) @Mclovin21 \(\text{3. Which equation represents a line with a slope}~ \color{red}{1/2}~\text{that passes through}~ \color{red}{(-2, 1)}\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 Now, if I compare the equation I wrote and the question, I see that: (1) The line passes through (-2,1),so, (-2,1) is the same as \((x_1,y_1)\). So, \(x_1 = -2\); \(y_1 = 1\). (2) The line has a slope 1/2, but, for the equation I wrote, the slope was m. So, \(m = \cfrac{1}{2}\). Thus, the equation of the required line will be: \(\large{(y-y_1) = m(x-x_1)}\) or, \(\large{(y-1) = \cfrac{1}{2}(x-(-2))}\) or, \(\large{(y-1) = \cfrac{1}{2}(x+2)}\) or, \(\large{(y-1)*2 = 1*(x+2)}\) or, \(\large{2y - 2 = x+2}\) or, \(\large{2y-x = 4}\) This is the require equation line. Observe this solution carefully and try solving your question. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

how did you get the *2 after (y-1) in the equation (y - 1) *2 = 1 * (x + 2)??

OpenStudy (vishweshshrimali5):

I actually multiplied 2 both sides, so that I could cancel out the 2 in denominator in RHS.